OpenStudy (anonymous):

Guys, kindly help.. calculus Find d/dx(2y/5x)

hartnn (hartnn):

@Flava Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) since you are differentiating with respect to x, y will be treated as constant(unless y is the function of x) so you can pull that out , giving (2y/5) d/dx (1/x) can u differentiate 1/x ?

OpenStudy (anonymous):

not sure, im still green in this. what's the answer

hartnn (hartnn):

1/x can be written as x^(-1) and differentiation of x^n is n x^{n-1} now can u differentiate x^(-1) ?? juat put n=-1 in the formula.

hartnn (hartnn):

*just

hartnn (hartnn):

@Flava we do not just give answers here, please read http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

im working it out, sorry you were too fast. I actually want to understand not just the answer,

hartnn (hartnn):

that is good. I will wait till you calculate d/dx (1/x)

OpenStudy (anonymous):

d/dx(1/x) = x^-1 how do you proceed from here

hartnn (hartnn):

thats not correct. 1/x = x^(-1) d/dx(1/x) = d/dx(x^(-1)) = ??

OpenStudy (anonymous):

=-x^2

hartnn (hartnn):

it will be -x^(-2) so d/dx (2y/5x) = 2y/5 * (-x^(-2)) or -2y/5x^2 ok?

OpenStudy (anonymous):

thanks so much, highly appreciated

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

hartnn - can we try another one... if 3x^2+2x^2y^3-5/4y^2=0, evaluate dy/dx when x=1/2 and y=1 dy/dx = 6x+4x3y^2-5y-2 then substiute x..?

OpenStudy (anonymous):

@Flava that last question uses implicit differentiation, once done then plug the values of x and y implicit differentiation is different to normal differentiation

OpenStudy (anonymous):

thanks for advis pls help, im not good on implict diff'. I understand the normal differentiation

OpenStudy (anonymous):

@Flava to tell you the truth I can solve it, but I find explaining implicit is very difficult, when you study it you will know what I mean, maybe @amistre64 or @UnkleRhaukus can help explaining it, or guide you elsewhere

OpenStudy (anonymous):

ok, how do you solve it. When i see the workings i will be able to see the difference from the normal differenciation

OpenStudy (unklerhaukus):

\[\qquad\qquad\qquad3x^2+2x^2y^3-\frac54y^2=0\]\[6x+4xy^3+6x^2y^2\frac{\text dy}{\text dx}-\frac52y\frac{\text dy}{\text dx}=0\]

OpenStudy (anonymous):

thanks Unkle

OpenStudy (unklerhaukus):

\[6x+4xy^3+\left(6x^2y^2-\frac52y\right)\frac{\text dy}{\text dx}=0\] \[6x+4xy^3=\left(\frac52y-6x^2y^2\right)\frac{\text dy}{\text dx}\] \[\frac{\text dy}{\text dx}=\frac{6x+4xy^3}{\left(\frac52y-6x^2y^2\right)}\]

OpenStudy (unklerhaukus):

now sub in \((1/2,1)\)

OpenStudy (amistre64):

the confusion and difficulty difference between explicit and implicit differentation stems from the way that a derivative is taught to begin with. We are simply not taught how vital the chain rule is to begin with.

OpenStudy (amistre64):

if we allow x and y to be functions of some unknown variable then we simply apply the chain rule. for example, D[4x^3] = 3*4x^2 * (x') it is only in using the information that the derivative is with respect to "x" that we can then simplify this to x' = 1

OpenStudy (unklerhaukus):

maybe Leibniz's notation is the most transparent here

OpenStudy (amistre64):

i agree; if we chain out a dx/dx its quite obvious what the behind the scenes workings are

OpenStudy (anonymous):

Unckle Thanks I was offline a bit. Aftr substituting I have got 5 as the answer Amsistre64 - i agree with you,most lecturers do not give clear disptinction between the two differentiation hence confising us

OpenStudy (anonymous):

Hartnn / Uncle kindly check for me if I've done this right... Find the gradient of the tangent drawn to the circle x^2+y^2=16 at the point where x=2 (current to 4s.f) My Soln: d/dx[x^2+y(x)^2=2x+2ydy/dx=0 2ydy/dx=-2x dy/dx = -x/y at x=2 -2/y

OpenStudy (unklerhaukus):

five is good fot the first question

OpenStudy (anonymous):

Uncle - thanks, what of the 2nd one

OpenStudy (unklerhaukus):

your working looks fine

OpenStudy (unklerhaukus):

there should be tow solutions right?

OpenStudy (anonymous):

what's the other soln?

OpenStudy (unklerhaukus):